# Stabilizing the Splits through Minimax Decision Trees

### v2

@@ Abstract
We prove oracle inequalities that cover both regression and classification, under mild marginal non-atomicity conditions. The bounds control the tree’s global excess risk by local worst-case impurities and yield fast convergence rates compared to CART. We extend the analysis to ensembles that subsample coordinates per node.
Empirically, (cyclic) MinimaxSplit trees and their forests improve on structured heterogeneity data such as EEG amplitude regression over fixed time horizons, seasonal air quality forecasting, and image denoising framed as non-parametric regression on spatial coordinates.
Keywords: CART; convergence rates; decision tree models; non-parametric regression
@@ 1 Introduction
Tree-based predictors are classical non-parametric learners, from early systems such as THAID to the CART formalization (Morgan and Messenger, 1973; Breiman et al., 1984), with surveys charting both breadth and limitations (Gordon and Olshen, 1984; Loh, 2014) and recent work revisiting large-scale aspects (Klusowski and Tian, 2024). A persistent and practically important pathology of single-tree greedy splitting is the end-cut preference (ECP): standard impurity/variance criteria favor splits near node boundaries, especially under weak or no signal. ECP has long been observed empirically (Morgan and Messenger, 1973; Torgo, 2001; Su et al., 2024; Wang, 2025) and is now rigorously characterized in simple settings, where noisy directions drive splits toward extremes (Breiman et al., 1984; Cattaneo et al., 2022).
For a single deep tree, ECP can be harmful: aggressive boundary splits create highly unbalanced children, amplify estimation noise, and can stall meaningful refinement along informative coordinates. However, Ishwaran (2015) argues that the same tendency can be beneficial in deep random forests: when a candidate split falls on a noise variable, an end-cut keeps one child large, preserving downstream sample size so later splits (possibly on strong variables) can recover the signal; even along strong variables, ECP-like behavior can be adaptive in regions with little local signal. Thus, while ECP degrades a single tree’s stability, it may aid an ensemble by lowering tree correlation and enabling recovery via aggregation. At the same time, the forest setting introduces its own costs: ECP interacts with candidate-set multiplicity and variable-usage bias, skewing split frequencies and importance metrics, and further reducing model interpretability. Existing methods to mitigate ECP include the delta splitting rule (Morgan and Messenger, 1973), regularization of node size (Torgo, 2001), and the Smooth Sigmoid Surrogate (Su et al., 2024). We refer to Wang (2025) for a more comprehensive review of the ECP.
@@
Our focus is on explicitly mitigating ECP at the algorithmic level for trees, while retaining the transparency that makes single decision trees interpretable. Beyond prediction, this design goal is interpretability: balanced, nondegenerate splits yield partitions whose geometry is easier to understand and communicate than the heterogeneous, boundary-driven partitions that arise under ECP in deep trees and ensembles.
In this paper, we revisit ECP empirically and theoretically, clarifying the cases in which standard CART-style criteria drift toward extreme cut-off points and how this differs between single trees and forests (Breiman et al., 1984; Morgan and Messenger, 1973; Torgo, 2001; Su et al., 2024; Ishwaran, 2015; Wang, 2025; Cattaneo et al., 2022). Then, we introduce a Minimax splitting rule with a cyclic coordinate schedule that systematically discourages degenerate child nodes, thereby mitigating ECP while preserving interpretability. As a consequence, our splitting rule admits a faster convergence guarantee compared to standard CART — we establish exponential error decay under mild conditions and relate our guarantees to existing theory (Chi et al., 2022; Syrgkanis and Zampetakis, 2020; Mazumder and Wang, 2024; Cattaneo et al., 2024). Conceptually, the analysis connects to partition-based martingale approximations (Simons, 1970; Zhang et al., 2024) and complements asymptotic and inferential developments for trees/forests (Klusowski and Tian, 2024; Chi et al., 2022; Syrgkanis and Zampetakis, 2020; Mazumder and Wang, 2024; Cattaneo et al., 2024), as well as recent scalable partitioning frameworks (Liu et al., 2023; Luo and Pratola, 2023; Luo et al., 2024) and complexity-theoretic perspectives on hierarchical splits (Blanc et al., 2020; O’Donnell et al., 2005).
Figure 1: The comparison of the scikit-learn DecisionTreeRegressor, our MinimaxSplit tree, and VarianceSplit tree on a sinusoidal target function (in red) $Y=f(X)+\varepsilon=\sin(4\pi X)+\varepsilon$, $X\buildrel\mathrm{law}\over{\sim}\mathrm{Unif}(0,1),$ and $\varepsilon\buildrel\mathrm{law}\over{\sim}\mathcal{N}(0,0.1)$, with maximum depth 4 (top row) and 5 (bottom row). We display 10 different model fits based on 10 different batches of training sets of size 100, and the average mean squared error (MSE) evaluated on a different training set across 10 batches.
@@ 2 Minimax decision trees for regression
The goal of this section is to construct MinimaxSplit decision tree algorithms that avoid ECP and attain exponential convergence. We first focus on the regression setting and we discuss the classification setting in Section 4. In CART-style regression, a decision tree greedily splits in order to minimize the sum of variances over the responses of the child nodes (Liu et al., 2023; Breiman et al., 1984); instead, we propose to greedily minimize the maximum variance among the child nodes. In Section 2.1, we rigorously introduce the algorithms (summarized in Figure 2). We study exponential rates for marginally atomless joint distributions (that may not arise from empirical measures $\mathbb{P}_{N}$) in Section 2.2. Section 2.3 builds upon this case, covering more generally the non-marginally atomless setting and, in particular, the empirical risk bounds. All proofs can be found in Appendix B.1. Throughout, we use the notation $d]=\{1,2,\dots,d\}$ for $d\in\mathbb{N}$. All splits we consider in this paper will be binary.
@@ 2.1 Algorithm formulation
Formally, suppose that a hyper-rectangle $A=[a_{1},b_{1})\times\dots\times[a_{d},b_{d})$ belongs to the partition $\pi_{k-1}$ and is splittable (if $A$ is unsplittable, we do not split $A$ and $A\in\pi_{\ell}$ for every $\ell\geqslant k$). In particular, $M_{k-1}(\omega)=\mathbb{E}[Y\mid\mathbf{X}\in A]$ if $\mathbf{X}(\omega)\in A$. To find a split of $A$ at depth $k$, the VarianceSplit algorithm looks for a dimension $j\in[d]$ and covariate $x_{j}\in[a_{j},b_{j})$ such that the remaining risk is the smallest after splitting $A$ at $x_{j}$ in covariate $j$. In other words, one seeks for
@@
$M_{k}(\omega)=\begin{cases}\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}<x_{j}]&\text{ if }\mathbf{X}(\omega)\in A_{L};\\ \mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}\geqslant x_{j}]&\text{ if }\mathbf{X}(\omega)\in A_{R}.\end{cases}$ (2)
@@
This leads to an associated process $\{M_{k}\}_{k\geqslant 0}$ to the joint distribution $(\mathbf{X},Y)$. [^2] The process $\{M_{k}\}_{k\geqslant 0}$ is a martingale, but we will not use this fact in the rest of the main text. To connect $M_{k}$ with the estimator $\hat{m}_{k}$ of the regression function at step $k$, we note that for any $\mathbf{x}\in\mathbb{R}^{d}$, there exists a unique $A_{\mathbf{x}}\in\pi_{k}$ containing $\mathbf{x}$, then $\hat{m}_{k}(\mathbf{x})=\mathbb{E}[Y\mid\mathbf{X}\in A_{\mathbf{x}}]$.
Figure 2: The comparison of the VarianceSplit, MinimaxSplit and cyclic MinimaxSplit trees. Here we use the notation: $A_{j,s}^{L}=A_{t}\cap\{x_{j}<s\}$$A_{j,s}^{R}=A_{t}\cap\{x_{j}\geqslant s\}$$\mathcal{S}_{j}(A_{t})$ are candidate thresholds (e.g., midpoints of sorted $x_{j}$ in $A_{t}$); $P_{t}(B)=\mathbb{P}(\mathbf{X}\in B\mid\mathbf{X}\in A_{t})$.
@@
The cyclic MinimaxSplit algorithm is a variation of the MinimaxSplit algorithm. Instead of optimizing over all covariates $j\in[d]$ in (4), we cycle through the $d$ covariates as the tree grows. That is, for $k\geqslant 0$, let $j=j_{k}=1+(k\leavevmode\nobreak\ (\text{mod }d))$. For all splittable sets $A\in\pi_{k}$, we split $A$ in the $j$-th coordinate to form its two descendants at depth $k+1$. The split location is then defined as
@@ 2.2 Risk bound
We note here a crucial difference from the statistics literature that our analysis of the MinimaxSplit rule first focuses on general joint distributions $(\mathbf{X},Y)$ on $\mathbb{R}^{d+1}$, where $\mathbf{X}\in\mathbb{R}^{d}$ and $Y\in\mathbb{R}$. [^3] This can be intuitively regarded as empirical measures with infinitely many samples. Empirical measures are a special case. The class of joint distributions we consider will satisfy the following marginally atomless property.
@@
We revisit (a special case of) the location regression model analyzed in Cattaneo et al. (2022). Let $(X,Y)$ be an empirical measure of $\mathrm{Unif}(0,1)\times\mathcal{N}(0,1)$ on $0,1]\times\mathbb{R}$ with $n$ samples. [^4] XX is not marginally atomless, but the example can always be modified by smoothing the data without affecting its validity. Suppose that we perform VarianceSplit for one step. Theorem 4.1 of Cattaneo et al. (2022) implies that, for $n$ large enough, the split lies between the first and $\sqrt{n}$-th order statistics with probability exceeding $1/6$. With high probability (meaning 1o(1)1-o(1) as $n\to\infty$), the right node has remaining risk at least (1o(1))n(1-o(1))n while the total risk at the parent node is at most $(1+o(1))n$. By a union bound, we see that with an arbitrary $\varepsilon>0$, for sufficiently large $n$,
@@
This example also explains why assumptions such as the sufficient impurity decay (SID) condition are favorable to ensure risk decay when using the splitting criterion (1) (Chi et al., 2022; Mazumder and Wang, 2024).
To see further intuition of ECP, we consider the event $\{\mathbf{X}\in A\}$ and sub-events $A_{L}=\left\{\mathbf{X}\in A,X_{j}<\hat{x}_{j}\right\}$ and $A_{R}=\left\{\mathbf{X}\in A,X_{j}\geqslant\hat{x}_{j}\right\}$. Without loss of generality, we may assume $\mathbb{P}(\mathbf{X}\in A)=1$ and apply the total variance formula to this partition, yielding
@@
From this, we see that when ECP occurs, $\mathbb{P}(A_{L})\approx 1$ or 0, the improvement is dominated by the first two terms of (12): $\mathrm{Var}(Y\mid A_{L})\mathbb{P}(A_{L})+\mathrm{Var}(Y\mid A_{R})\mathbb{P}(A_{R})=V_{\text{left}}(j)+V_{\text{right}}(j)$. Therefore, when ECP occurs, the sum of variances does not necessarily decay fast enough for the VarianceSplit.
However, for (cyclic) MinimaxSplit, $\max\{V_{\text{left}}(j),V_{\text{right}}(j)\}$ always displays exponential decay (11), hence “avoiding” the ECP phenomenon. In other words, the fast decay of the sum of variances is not easily guaranteed from the VarianceSplit construction due to ECP; but the exponential decay of the maximum variance can be assured by the MinimaxSplit construction, as we show in Theorem 2.
@@
where $\mathbf{x}=(x_{1},\dots,x_{d})$. For $g\in\mathcal{G}$, define $\left\lVert g\right\rVert_{\mathrm{TV}}$ as the infimum of $\sum_{i=1}^{d}\left\lVert g_{i}\right\rVert_{\mathrm{TV}}$ over all such additive representations of $g$, where $\left\lVert\cdot\right\rVert_{\mathrm{TV}}$ denotes the total variation of a univariate function of bounded variation.
@@
Theorem 2 does not hold in general for $d\geqslant 2$ if we replace the cyclic MinimaxSplit algorithm by the MinimaxSplit algorithm, a phenomenon we illustrate in Section 2.4 below. Nevertheless, the MinimaxSplit algorithm often achieves desirable risk decay in high-dimensional settings in which cyclic MinimaxSplit is inefficient; see Appendix D.2.
@@
Theorems 6 and 7 below enjoy the same specialized bounds.
@@
$\mathbb{E}\bigl{[}(Y-M_{k})^{2}\bigr{]}=\sum_{i=k}^{\infty}\mathbb{E}\bigl{[}(M_{i+1}-M_{i})^{2}\bigr{]}\quad\text{ and }\quad\mathbb{E}\bigl{[}(M_{k}-M_{k+1})^{2}\bigr{]}=\sum_{A\in\pi_{k}}\sum_{\begin{subarray}{c}A^{\prime}\in\pi_{k+1}\\ A^{\prime}\subset A\end{subarray}}\mathbb{P}(A^{\prime})\bigl{(}\mu_{A^{\prime}}-\mu_{A}\bigr{)}^{2}.$ (18)
In other words, convergence rates reduce to controlling the one-step increments created by each split. The VarianceSplit rule provides no uniform control: under ECP, a child can have arbitrarily small mass, so large within-node fluctuations can be confined to microscopic slivers, and the increment need not contract at a fixed rate. Our MinimaxSplit instead minimizes $\max\{\mathbb{P}(A_{L})\mathrm{Var}(Y\mid A_{L}),\,\mathbb{P}(A_{R})\mathrm{Var}(Y\mid A_{R})\}$, ensuring that at every refinement the worst child risk decreases by a constant factor. Summed along the tree, this yields an exponential contraction of $\mathbb{E}[(M_{k+1}-M_{k})^{2}]$ and hence of $\mathbb{E}[(Y-M_{k})^{2}]$. The cyclic schedule strengthens this guarantee by forcing each coordinate to be refined regularly, preventing all contraction from being spent on a single axis and avoiding symmetry-breaking failures; see Section 2.4.
@@
Consider the case where $d=1$, $X\buildrel\mathrm{law}\over{\sim}\mathrm{Unif}(0,1)$, and $Y=\sin(2\pi 2^{p}X)$. Figure 4 below plots the remaining risk $\mathbb{E}[(Y-M_{k})^{2}],\leavevmode\nobreak\ 1\leqslant k\leqslant K$ with power parameters $p=1,2,3,4$ and maximum tree depth $K=12$. Due to the symmetry of the sine function, the first few splits of the MinimaxSplit do not reduce the remaining risk, since the splitting points are points of symmetry. However, as soon as the splitting breaks through the symmetries, the remaining risk decays exponentially to zero. The VarianceSplit algorithm makes mild progress at every step, but is quickly surpassed by MinimaxSplit once the latter breaks through the symmetries. The reason is that due to ECP, the splitting points are near the boundary of the nodes, leading to unbalanced nodes. In other words, while the MinimaxSplit algorithm may not reduce the risk as steadily as VarianceSplit, the steps with little risk reduction often implicitly overcome certain hard constraints inherent to the regression problem (such as the symmetries of the sine function), paving the way for a much faster risk decay in later stages. Note that the occurrence of minimal risk deduction from the first few steps of MinimaxSplit does not contradict the risk bound from Theorem 2, because the upper bound already incorporates the total variation of the signal function.
Figure 4: Comparison of the mean squared errors for the MinimaxSplit and VarianceSplit regimes, in the setting where $X\buildrel\mathrm{law}\over{\sim}\mathrm{Unif}(0,1)$ and $Y=\sin(2\pi 2^{p}X)$ with $p=1,2,3,4$ and maximum tree depth $12$.

### v3

@@ Abstract
We prove oracle inequalities that cover both regression and classification, under mild marginal non-atomicity conditions. The bounds control the tree’s global excess risk by local worst-case impurities and yield fast convergence rates compared to CART. We extend the analysis to a random-dimension forest variant that subsamples coordinates per node.
Empirically, (cyclic) MinimaxSplit trees and their forests improve over baselines on structured heterogeneous data such as EEG amplitude regression over fixed time horizons and image denoising, framed as non-parametric regression on spatial coordinates.
Keywords: CART; convergence rates; decision tree models; non-parametric regression; classification
@@ 1 Introduction
Tree-based predictors are classical non-parametric learners, from early systems such as THAID to the CART formalization (Morgan and Messenger, 1973; Breiman et al., 1984), with surveys charting both breadth and limitations (Gordon and Olshen, 1984; Loh, 2014) and recent work revisiting large-scale aspects (Klusowski and Tian, 2024). A persistent and practically important pathology of single-tree greedy splitting is the end-cut preference (ECP): standard impurity/variance criteria favor splits near node boundaries, especially under weak or no signal. ECP has long been observed empirically (Morgan and Messenger, 1973; Torgo, 2001; Su et al., 2024; Wang, 2025) and is now rigorously characterized in simple settings, where noisy directions drive splits toward extremes (Breiman et al., 1984; Cattaneo et al., 2022). A simple example of ECP is given by Figure 1 below; see also Figure 4.
For a single deep tree, ECP can be harmful: aggressive boundary splits create highly unbalanced children, amplify estimation noise, and can stall meaningful refinement along informative coordinates. However, Ishwaran (2015) argues that the same tendency can be beneficial in deep random forests: when a candidate split falls on a noise variable, an end-cut keeps one child large, preserving downstream sample size so later splits (possibly on strong variables) can recover the signal; even along strong variables, ECP-like behavior can be adaptive in regions with little local signal. Thus, while ECP degrades a single tree’s stability, it may aid an ensemble by lowering tree correlation and enabling recovery via aggregation. At the same time, the forest setting introduces its own costs: ECP interacts with candidate-set multiplicity and variable-usage bias, skewing split frequencies and importance metrics, and further reducing model stability. Existing methods to mitigate ECP include the delta splitting rule (Morgan and Messenger, 1973), regularization of node size (Torgo, 2001), and the Smooth Sigmoid Surrogate (Su et al., 2024). We refer to Wang (2025) for a more comprehensive review of the ECP.
@@
Our focus is on explicitly mitigating ECP at the algorithmic level for trees, while retaining the transparency that makes single decision trees stable. Beyond prediction, this design goal is to generate balanced, nondegenerate splits that yield partitions whose geometry is easier to understand and communicate than the heterogeneous, boundary-driven partitions that arise under ECP in deep trees and ensembles.
In this paper, we revisit ECP empirically and theoretically, clarifying the cases in which standard CART-style criteria drift toward extreme cut-off points and how this differs between single trees and forests (Breiman et al., 1984; Morgan and Messenger, 1973; Torgo, 2001; Su et al., 2024; Ishwaran, 2015; Wang, 2025; Cattaneo et al., 2022). Then, we introduce a Minimax splitting rule with a cyclic coordinate schedule that systematically discourages degenerate child nodes, thereby mitigating ECP while preserving stability. As a consequence, our splitting rule admits a faster convergence guarantee compared to standard CART — we establish exponential error decay under mild conditions and relate our guarantees to existing theory (Chi et al., 2022; Syrgkanis and Zampetakis, 2020; Wager and Walther, 2015; Mazumder and Wang, 2024; Cattaneo et al., 2024). Conceptually, the analysis connects to partition-based martingale approximations (Simons, 1970; Zhang et al., 2024) and complements asymptotic and inferential developments for trees/forests (Klusowski and Tian, 2024; Chi et al., 2022; Syrgkanis and Zampetakis, 2020; Wager and Walther, 2015; Mazumder and Wang, 2024; Cattaneo et al., 2024), as well as recent scalable partitioning frameworks (Liu et al., 2023; Luo and Pratola, 2023; Luo et al., 2024; Liu and Ma, 2024) and complexity-theoretic perspectives on hierarchical splits (Blanc et al., 2020; O’Donnell et al., 2005).
Figure 1: The comparison of the scikit-learn DecisionTreeRegressor, our MinimaxSplit tree, and VarianceSplit tree on a sinusoidal target function (in red) $Y=f(X)+\varepsilon=\sin(4\pi X)+\varepsilon$, $X\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\mathrm{law}}}\mathrm{Unif}(0,1),$ and $\varepsilon\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\mathrm{law}}}\mathcal{N}(0,0.1)$, with maximum depth 4 (top row) and 5 (bottom row). We display 10 different model fits based on 10 different batches of training sets of size 100, and the average mean squared error (MSE) evaluated on a different training set across 10 batches.
@@ 2 Minimax decision trees for regression
The goal of this section is to construct (cyclic) MinimaxSplit decision tree algorithms that avoid ECP and attain exponential convergence. We first focus on the regression setting and we discuss the classification setting in Section 4. In short, in CART-style regression, a decision tree greedily splits in order to minimize the sum of variances over the responses of the child nodes (Liu et al., 2023; Breiman et al., 1984); instead, we propose to greedily minimize the maximum variance among the child nodes. The roadmap of this section is as follows.
  • In Section 2.1, we rigorously introduce the algorithms (summarized in Figure 2). Crucially, we distinguish two distinct algorithms, where we may optimize over all dimensions at each split (the MinimaxSplit algorithm) or we may pre-select a cyclic schedule of dimensions for each depth (the cyclic MinimaxSplit algorithm).
  • In Section 2.2, we study exponential rates for marginally atomless joint distributions (that may not arise from empirical measures $\mathbb{P}_{N}$) for the cyclic MinimaxSplit algorithm.
  • Section 2.3 builds upon the marginally atomless case, covering more generally the non-marginally atomless setting and, in particular, the empirical risk bounds for the cyclic MinimaxSplit algorithm.
  • In Section 2.4, we discuss why there are no general analogous results for the MinimaxSplit algorithm, even though it is shown to perform well empirically even in high-dimensional settings (Appendix E.3). Intuitively, the MinimaxSplit algorithm may fail to break symmetry if strong symmetry is imposed in the underlying samples.
All proofs can be found in Appendix B.1. Throughout, we use the notation $d]=\{1,2,\dots,d\}$ for $d\in\mathbb{N}$. All splits we consider in this paper will be binary.
@@ 2.1 Algorithm formulation
Formally, suppose that a hyper-rectangle $A=[a_{1},b_{1})\times\dots\times[a_{d},b_{d})$ belongs to the partition $\pi_{k-1}$ and is splittable (if $A$ is unsplittable, we do not split $A$ and $A\in\pi_{\ell}$ for every $\ell\geqslant k$). In particular, $M_{k-1}(\omega)=\mathbb{E}[Y\mid\mathbf{X}\in A]$ if $\mathbf{X}(\omega)\in A$. To find a split of $A$ at depth $k$, the VarianceSplit algorithm looks for a dimension $j\in[d]$ and covariate $x_{j}\in[a_{j},b_{j})$ such that the remaining risk is the smallest after splitting $A$ at $x_{j}$ in covariate $j$. In other words, one looks for
@@
$M_{k}(\omega)=\begin{cases}\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}<\hat{x}_{j}]&\text{ if }\mathbf{X}(\omega)\in A_{L};\\ \mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}\geqslant\hat{x}_{j}]&\text{ if }\mathbf{X}(\omega)\in A_{R}.\end{cases}$ (2)
@@
This leads to an associated process $\{M_{k}\}_{k\geqslant 0}$ with the joint distribution $(\mathbf{X},Y)$. [^2] The process $\{M_{k}\}_{k\geqslant 0}$ is a martingale, but we will not use this fact in the rest of the main text. To connect $M_{k}$ with the estimator $\hat{m}_{k}$ of the regression function at step $k$, we note that for any $\mathbf{x}\in\mathbb{R}^{d}$, there exists a unique $A_{\mathbf{x}}\in\pi_{k}$ containing $\mathbf{x}$, and then $\hat{m}_{k}(\mathbf{x})=\mathbb{E}[Y\mid\mathbf{X}\in A_{\mathbf{x}}]$.
Figure 2: The comparison of the VarianceSplit, MinimaxSplit and cyclic MinimaxSplit trees. Here we use the notation: $A_{j,s}^{L}=A_{t}\cap\{x_{j}<s\}$, $A_{j,s}^{R}=A_{t}\cap\{x_{j}\geqslant s\}$; $\mathcal{S}_{j}(A_{t})$ are candidate thresholds (e.g., midpoints of sorted $x_{j}$ in $A_{t}$); $P_{t}(B)=\mathbb{P}(\mathbf{X}\in B\mid\mathbf{X}\in A_{t})$.
@@
The cyclic MinimaxSplit algorithm is a variation of the MinimaxSplit algorithm. Instead of optimizing over all covariates $j\in[d]$ in (4), we cycle through the $d$ covariates as the tree grows. That is, for $k\geqslant 0$, let $j=j_{k}=1+(k\penalty 10000\ (\text{mod }d))$. For all splittable sets $A\in\pi_{k}$, we split $A$ in the $j$-th coordinate to form its two descendants at depth $k+1$. The split location is then defined as
@@ 2.2 Risk bound
We note here that a crucial difference from the statistics literature is that our analysis of the MinimaxSplit rule first focuses on general joint distributions $(\mathbf{X},Y)$ on $\mathbb{R}^{d+1}$, where $\mathbf{X}\in\mathbb{R}^{d}$ and $Y\in\mathbb{R}$. [^3] This can be intuitively regarded as empirical measures with infinitely many samples. Empirical measures are a special case. Note that this does not mean that our analysis is restricted to the noiseless case; see the setting in Section 2.3. The class of joint distributions we consider will satisfy the following marginally atomless property.
@@
Remark 1.
The marginally atomless property of $\mathbf{X}$ is not required for the monotonicity properties of the maps $\varphi_{L}$ and $\varphi_{R}$, but it is necessary for the continuity. The continuity is essential for an exact $1/2$ risk decay bound (see below), but this argument will be relaxed later in the derivation of Theorem 7.
@@
We revisit (a special case of) the location regression model analyzed in Cattaneo et al. (2022). Let $(X,Y)$ follow the law of an empirical measure of $\mathrm{Unif}(0,1)\times\mathcal{N}(0,1)$ on $0,1]\times\mathbb{R}$ with $n$ samples. [^4] XX is not marginally atomless, but the example can always be modified by smoothing the data without affecting its validity. Suppose that we perform VarianceSplit for one step. Theorem 4.1 of Cattaneo et al. (2022) implies that, for $n$ large enough, the split lies between the first and $\sqrt{n}$-th order statistics with probability exceeding $1/6$. With high probability (meaning 1o(1)1-o(1) as $n\to\infty$), the right node has remaining risk at least (1o(1))(1-o(1)) while the total risk at the parent node is at most $(1+o(1))$. By a union bound, we see that with an arbitrary $\varepsilon>0$, for sufficiently large $n$,
@@
This example also explains why assumptions such as the sufficient impurity decay (SID) condition are favorable to ensure risk decay when using the splitting criterion (1) (Chi et al., 2022; Mazumder and Wang, 2024). Roughly speaking, the SID condition assumes that the best population split removes at least a positive fraction of the within-node variance, leading to a lower bound on local CART progress and eventually an exponential convergence rate in depth for CART and random forests.
To gain further intuition about ECP, we consider the event $\{\mathbf{X}\in A\}$ and sub-events $A_{L}=\left\{\mathbf{X}\in A,X_{j}<\hat{x}_{j}\right\}$ and $A_{R}=\left\{\mathbf{X}\in A,X_{j}\geqslant\hat{x}_{j}\right\}$. Without loss of generality, we may assume $\mathbb{P}(\mathbf{X}\in A)=1$ and apply the total variance formula to this partition, yielding
@@
From this, we see that when ECP occurs, $\mathbb{P}(A_{L})\approx 1$ or 0, the total remaining risk is dominated by the first two terms of (12): $\mathrm{Var}(Y\mid A_{L})\mathbb{P}(A_{L})+\mathrm{Var}(Y\mid A_{R})\mathbb{P}(A_{R})=V_{\text{left}}(j)+V_{\text{right}}(j)$. Therefore, when ECP occurs, the sum of variances does not necessarily decay fast enough under VarianceSplit.
@@
where $\mathbf{x}=(x_{1},\dots,x_{d})$. For $g\in\mathcal{G}$, define $\left\lVert g\right\rVert_{\mathrm{TV}}$ as the infimum of $\sum_{i=1}^{d}\left\lVert g_{i}\right\rVert_{\mathrm{TV}}$ over all such additive representations of $g$, where $\left\lVert\cdot\right\rVert_{\mathrm{TV}}$ denotes the total variation of a univariate function of bounded variation. Note that this is the same type of approximation target used by Klusowski and Tian (2024); Mazumder and Wang (2024) for CART/C4.5: their oracle inequality is also stated as an infimum over the additive function class $\mathcal{G}$, with a penalty depending on the additive total-variation norm.
However, for (cyclic) MinimaxSplit, $\max\{V_{\text{left}}(j),V_{\text{right}}(j)\}$ always displays exponential decay (11), hence “avoiding” the ECP phenomenon. In other words, the fast decay of the sum of variances is not easily guaranteed from the VarianceSplit construction due to ECP; but the exponential decay of the maximum variance can be assured by the MinimaxSplit construction, as we show in Theorem 3. The MinimaxSplit construction approximately equalizes the two child risks, which implies that the larger child risk is at most one half of the parent risk. This local worst-child contraction is the basic mechanism behind the global error bound below.
@@
Theorem 3 does not hold in general for $d\geqslant 2$ if we replace the cyclic MinimaxSplit algorithm by the MinimaxSplit algorithm, a phenomenon we illustrate in Section 2.4 below. Nevertheless, the MinimaxSplit algorithm often achieves desirable risk decay in high-dimensional settings in which cyclic MinimaxSplit is inefficient; see Appendix E.3.
@@
Theorems 7 and 8 below enjoy the same specialized bounds.
@@
$\mathbb{E}\bigl[(Y-M_{k})^{2}\bigr]=\sum_{i=k}^{\infty}\mathbb{E}\bigl[(M_{i+1}-M_{i})^{2}\bigr]\quad\text{ and }\quad\mathbb{E}\bigl[(M_{k}-M_{k+1})^{2}\bigr]=\sum_{A\in\pi_{k}}\sum_{\begin{subarray}{c}A^{\prime}\in\pi_{k+1}\\ A^{\prime}\subset A\end{subarray}}\mathbb{P}(A^{\prime})\bigl(\mu_{A^{\prime}}-\mu_{A}\bigr)^{2},$ (18)
where we have used that $Y$ is $\sigma(X)$-measurable. In other words, convergence rates reduce to controlling the one-step increments created by each split. The VarianceSplit rule provides no uniform control: under ECP, a child can have arbitrarily small mass, so large within-node fluctuations can be confined to microscopic slivers, and the increment need not contract at a fixed rate. Our MinimaxSplit instead minimizes $\max\{\mathbb{P}(A_{L})\mathrm{Var}(Y\mid A_{L}),\,\mathbb{P}(A_{R})\mathrm{Var}(Y\mid A_{R})\}$, ensuring that at every refinement the worst child risk decreases by a constant factor. Summed along the tree, this yields an exponential contraction of $\mathbb{E}[(M_{k+1}-M_{k})^{2}]$ and hence of $\mathbb{E}[(Y-M_{k})^{2}]$. The cyclic schedule strengthens this guarantee by forcing each coordinate to be refined regularly, preventing all contraction from being spent on a single axis and avoiding symmetry-breaking failures; see Section 2.4.
@@
Consider the case where $d=1$, $X\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\mathrm{law}}}\mathrm{Unif}(0,1)$, and $Y=\sin(2\pi 2^{p}X)$. Figure 4 below plots the remaining risk $\mathbb{E}[(Y-M_{k})^{2}],\penalty 10000\ 1\leqslant k\leqslant K$ with power parameters $p=1,2,3,4$ and maximum tree depth $K=12$. Due to the symmetry of the sine function, the first few splits of the MinimaxSplit do not reduce the remaining risk, since the splitting points are points of symmetry. However, as soon as the splitting breaks through the symmetries, the remaining risk decays exponentially to zero. The VarianceSplit algorithm makes mild progress at every step, but is quickly surpassed by MinimaxSplit once the latter breaks through the symmetries. The reason is that due to ECP, the splitting points are near the boundary of the nodes, leading to unbalanced nodes. In other words, while the MinimaxSplit algorithm may not reduce the risk as steadily as VarianceSplit, the steps with little risk reduction often implicitly overcome certain hard constraints inherent to the regression problem (such as the symmetries of the sine function), paving the way for a much faster risk decay in later stages. Note that the occurrence of minimal risk reduction from the first few steps of MinimaxSplit does not contradict the risk bound from Theorem 3, because the upper bound already incorporates the total variation of the signal function.
Figure 4: Comparison of the mean squared errors for the MinimaxSplit and VarianceSplit regimes, in the setting where $X\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\mathrm{law}}}\mathrm{Unif}(0,1)$ and $Y=\sin(2\pi 2^{p}X)$ with $p=1,2,3,4$ and maximum tree depth $12$.